# Sim-to-Real Transfer of Deep Reinforcement Learning Agents for Online Coverage Path Planning

### v2

@@ Introduction
Action space. We consider a differential drive wheeled robot, which is controlled by two separately driven wheels. The agent predicts the linear and angular velocities $v$ and $\omega$ for the robot. In contrast to the cited work, we convert them to angular velocities $\omega_{R}$ and $\omega_{L}$ for the right and left wheels,
@@
Reward function. In addition to a coverage reward term, Jonnarth, Zhao, and Felsberg (2024) found it necessary to include a total variation (TV) reward that penalizes variations in the coverage map, such as excessive boundaries. By reducing the total variation, the agent learns to cover the area in a more compact manner, reducing leftover holes in the coverage map, and thus, allows for complete coverage. In total, four reward terms are used: a coverage reward $R_{\mathrm{area}}$, a TV reward $R_{\mathrm{TV}}$, a collision reward $R_{\mathrm{coll}}$, and a constant reward $R_{\mathrm{const}}$. The coverage reward is written as
@@
where $\lambda_{\mathrm{area}}$ is a scaling factor, $A_{\mathrm{new}}$ is the newly covered area, $v_{\mathrm{max}}$ is the maximum speed of the robot, and $\Delta t$ is the time step size. The denominator is the largest possible area that can be covered in a time step and ensures that the reward is normalized to $[0,\lambda_{\mathrm{area}}]$. The TV reward is written as
$R_{\mathrm{TV}}(t)=-\lambda_{\mathrm{TV}}\frac{V(C_{t})-V(C_{t-1})}{2v_{\mathrm{max}}\Delta t},$ (3)
with scaling factor $\lambda_{\mathrm{TV}}$, where $C_{t}$ is the coverage map at time $t$, and $V$ is the total variation
$V(x)=\sum_{i,j}\sqrt{|x_{i+1,j}-x_{i,j}|^{2}+|x_{i,j+2}-x_{i,j}|^{2}}.$ (4)
Finally, the agent is penalized by a negative obstacle collision reward $R_{\mathrm{coll}}$ on impact, and a constant negative reward $R_{\mathrm{const}}$ is given each time step. Each episode is terminated when a goal coverage has been reached, or after $\tau=1000$ consecutive steps without new ground being covered.
Model architecture. To process the multi-scale maps together with the sensor data, a scale-grouped convolutional neural network (SGCNN) is used, which has shown to be superior to a standard CNN or MLP (Jonnarth, Zhao, and Felsberg 2024). It consists of a map feature extractor $g_{m}$, a sensor feature extractor $g_{s}$, and a fusing module $g_{f}$. The map feature extractor groups the maps by scale and processes them separately using grouped convolutions, while the sensor feature extractor and fusing module consist of fully connected layers. The control signal predictions are given by
$(v,\omega)=g_{f}(g_{m}(M_{c},M_{o},M_{f}),g_{s}(S)),$ (5)
where $M_{c}$, $M_{o}$, and $M_{f}$ are the multi-scale coverage, obstacle, and frontier maps respectively, and $S$ is a vector of the observed normalized lidar distance measurements.
Transferring CPP Agents to the Real World
The approach described in the previous section produces CPP agents with state-of-the-art performance in simulation. Obviously, one could train the same agent directly in the real environment. However, this would require training with manual interaction in the order of weeks to months. In this section, we describe our proposed approach, which allows to transfer the model trained in simulation to the real robot.
Learning CPP on the Real System
@@
Even if we will not train the agent from scratch in the real setting, we need to be able to fine-tune the transferred system by RL. In common RL-libraries, such as stable-baselines3 (Raffin et al. 2021), tianshou (Weng et al. 2022), and spinningup (Achiam 2018), the data collection and model updates are performed serially. While this is feasible in simulation and turn-based environments where the environment can be paused during model updates and time can be advanced instantly to the next state after action selection, it is not practical for real-time robotic applications where the agent is trained online.
@@
In this case, we need to wait after action selection to observe the next state, while the execution remains idle. After the environment interaction, a batch is sampled from the replay buffer, gradients are computed, and the model is updated. However, this takes time, which is not negligible, especially on low-performance edge devices and embedded systems. During the gradient update step, the robot keeps following its previous action. As a result, the agent ends up in a state which deviates from the recorded one used for action selection. Following previous work (Haarnoja et al. 2019; Yuan and Mahmood 2022), we perform the model updates in parallel with the data collection, utilizing computational resources which would otherwise remain idle while waiting for the next state.
Figure 1: During online training we perform environment interactions (top) and model training (bottom) in parallel.
@@
The training process can be divided into four main computational steps, where the respective measured times are given in Table 1:
(1) Action selection:
@@
Given the current observation, we sample an action from the policy. This corresponds to a forward pass of the policy network.
(2) State selection:
@@
After sampling an action, the control signals are sent to the robot platform, and the new state is observed. In simulation, this would occur instantly, while in the real world we need to let time pass to observe how the action affected the environment. After state selection, the reward is computed, and the action, previous state, new state, and reward are added to the replay buffer.
(3) Batch sampling:
A training batch is sampled from the replay buffer.
(4) Model update:
@@
Gradients are computed based on the training batch, and the model weights are updated. This is the most computationally intensive part.
@@
Table 1: Measured times of the RL-step on the real system. Timing for (2) is application dependent, and chosen based on the optimal training step size. The overhead effectively becomes part of (1) resulting in an action delay of 50 ms.
Since the action selection and batch sampling were fast compared to the state selection and gradient computations, we only use two threads to reduce overhead and complexity. Our computational graph is shown in Figure 1, which consists of an environment interaction thread (top) and a model update thread (bottom). Since both threads interact with the replay buffer, we use a mutex to avoid conflicts. Luckily, the the environment interaction thread adds an entry at the end of the cycle, while the model update thread samples from it early, so the there is a low chance that they block each other.
Since the model update thread modifies the weight tensors during a large portion of the runtime, i.e. both during the backward pass to update the gradients and during the optimizer step to update the parameters, a mutex is not feasible as it would block execution too often. Instead, we keep a copy of the model, which is only accessed by the model update thread. The weights are synchronized when both threads have finished.
@@
Smoothening the Sim-to-Real Gap
With the approach in the previous section, we could directly move to a real setting in-the-wild. However, this would make training and fine-tuning cumbersome as manual interaction would be required to supervise the training process. In order to fully automatically (continue to) train the system, we propose to use a semi-virtual setup, see Figure 2. In this setup, we use the real robot with its kinematics and dynamics, but simulate its lidar sensor and obstacles, and localize the robot using a positioning system. The environment, including obstacles and coverage, e.g. from mowing, is visualized by a projection onto the ground.
The main purpose of using such an environment is to evaluate, in a controlled manner, the ability of the learning algorithm to generalize to a real robot in a real-time setting. Furthermore, the considered RL approach is robust to sensing and localization noise, and can be extended to the case of unknown pose, e.g. by estimating the pose using an off-the-shelf SLAM method (Jonnarth, Zhao, and Felsberg 2024). Thus, using our semi-virtual environment, we can draw conclusions about how we expect similar policies to generalize to fully realistic settings.
Figure 2: Picture of the robot having covered an environment with four obstacles.
@@
To further smoothen the sim-to-real gap, we improve the fidelity of the simulator by taking into consideration the latencies induced by inertia and action delay. We measure the maximum linear acceleration, maximum angular acceleration, and action delay of the real system, and include these aspects in the simulated kinematics and dynamics.
To account for higher-order Markovian dynamics, we include information from previous time steps in the observation space. While the common approach is to stack several previous observations (Haarnoja et al. 2019; Mnih et al. 2013), it is less feasible in our setting for multiple reasons. (1) Since the pose is embedded in the egocentric maps, rapid rotations significantly alter the observed state, making it difficult to learn the dynamics. (2) It significantly increases the model size and processing time, which are critical aspects for real-time applications. (3) It severely limits the replay buffer size, as the map observations are fairly large. Instead, we use a history of the previous actions (Haarnoja et al. 2019; Mnih et al. 2013), which is lightweight and avoids the listed problems, and should be sufficient for learning the dynamics of the robot. A benefit of using action observations instead of velocity estimates is that the estimates can be highly noisy. Furthermore, using action observations, the agent can learn to estimate the velocity if necessary.
Optimal Strategy for Going Sim-to-Real
@@
    How does the model with fine-tuning on the real robot perform in comparison to (2) and dependent on the number of training steps?
@@
Experiments
In this section, we present our experimental results for transferring a state-of-the-art CPP RL policy to a semi-virtual environment.
Implementation Details

### v3

@@ III-D Action Space
In the real-world setting, we use a differential drive wheeled robot, which is controlled by two separately driven wheels. The agent predicts the linear and angular velocities $v$ and $\omega$ for the robot, which are converted to angular velocities $\omega_{R}$ and $\omega_{L}$ for the right and left wheels,
@@
A continuous action space is of course not the only choice, but it is more realistic than a discrete one. There are many reasons why we chose continuous actions. (1) This allows us to model a continuous pose that is not constrained by the grid discretization, and thus not constraining the path space. (2) The agent can adapt and optimize its path for a specific kinematic model. In our experiments in Section V we find that our approach works well in continuous settings, which hints that it may also work in other continuous control spaces for specific kinematic models. The same conclusion could not be drawn if a discrete action space was used. Our specific choice of linear and angular velocities applies directly to a wide variety of robots, e.g. differential drive systems and the Ackermann kinematic model (by using a constraint on the angular velocity). This makes it more suitable for sim-to-real transfer compared to a discrete action space. (3) Compared to grid-based discrete actions, a continuous action space introduces additional sources for error, both regarding dynamics and localization, which are more realistic.
III-E Reward Function
As the goal is to cover the free space, a reward based on covering new ground is required. Similar to Chaplot et al. [43] and Chen et al. [41], we define a reward term $R_{\mathrm{area}}$ based on the newly covered area $A_{\mathrm{new}}$ in the current time step that was not covered previously. To control the scale of this reward signal, we first normalize it to $[0,1]$ based on the maximum area that can be covered in each time step, which is related to the maximum speed $v_{\mathrm{max}}$. We subsequently multiply it with the scale factor $\lambda_{\mathrm{area}}$, which is the maximum possible reward in each step, resulting in the reward
@@ III-E Reward Function
where $r$ is the agent radius and $\Delta t$ is the time step size. See Fig. 3 for an illustration.
Figure 3: An illustration of the area reward $R_{\mathrm{area}}$, which is based on the maximum possible area that can be covered in each time step.
Refer to caption Refer to caption
Figure 4: The total variation reward term improves path quality. The figure shows the path (yellow line) without TV reward (left), and with TV reward (right), including the covered region (green), lidar rays (blue lines) and frontier points (magenta).
By only maximizing the coverage reward in (2), the agent is discouraged from overlapping its previous path, as this reduces the reward in the short term. This leads to holes or stripes in the coverage, which we observed in our experiments, see Fig. 4. These leftover parts can be costly to cover later on for reaching complete coverage. Covering the thin stripes afterward only yields a minor reward, resulting in slow convergence towards an optimal path. To reduce the leftover parts, we propose a reward term based on minimizing the total variation (TV) of the coverage map. Minimizing the TV corresponds to reducing the boundary of the coverage map, and thus leads to fewer holes. Given a 2D signal $x$, the discrete isotropic total variation, which has been used for image denoising [44], is expressed as
$V(x)=\sum_{i,j}\sqrt{|x_{i+1,j}-x_{i,j}|^{2}+|x_{i,j+1}-x_{i,j}|^{2}}.$ (3)
We consider two variants of the TV reward term, a global and an incremental. For the global TV reward $R_{\mathrm{TV}}^{\mathrm{G}}$, the agent is given a reward based on the global coverage map $C_{t}$ at time step $t$. To avoid an unbounded TV for large environments, it is scaled by the square root of the covered area $A_{\mathrm{covered}}$, as this results in a constant TV for a given shape of the coverage map, independent of scale. The incremental TV reward $R_{\mathrm{TV}}^{\mathrm{I}}$ is based on the difference in TV between the current and the previous time step. A positive reward is given if the TV is decreased, and vice versa. The incremental reward is scaled by the maximum possible increase in TV in a time step, which is twice the traveled distance. The global and incremental rewards are respectively given by
$R_{\mathrm{TV}}^{\mathrm{G}}(t)$ $=-\lambda_{\mathrm{TV}}^{\mathrm{G}}\frac{V(C_{t})}{\sqrt{A_{\mathrm{covered}}}},$ (4)
$R_{\mathrm{TV}}^{\mathrm{I}}(t)$ $=-\lambda_{\mathrm{TV}}^{\mathrm{I}}\frac{V(C_{t})-V(C_{t-1})}{2v_{\mathrm{max}}\Delta t},$ (5)
where $\lambda_{\mathrm{TV}}^{\mathrm{G}}$ and $\lambda_{\mathrm{TV}}^{\mathrm{I}}$ are reward scaling parameters to make sure that $|R_{\mathrm{TV}}|<|R_{\mathrm{area}}|$ on average. Otherwise, the optimal behaviour is simply to stand still.
To avoid obstacle collisions, a negative reward $R_{\mathrm{coll}}$ is given each time the agent collides with an obstacle. Finally, a small constant negative reward $R_{\mathrm{const}}$ is given in each time step to encourage fast execution. Thus, our final reward function reads
$R=R_{\mathrm{area}}+R_{\mathrm{TV}}^{\mathrm{G}}+R_{\mathrm{TV}}^{\mathrm{I}}+R_{\mathrm{coll}}+R_{\mathrm{const}}.$ (6)
During training, each episode is terminated when the agent reaches a pre-defined goal coverage, or when it has not covered any new space in $\tau$ consecutive time steps.
III-F Agent Architecture
Due to the multi-modal nature of the observation space, we use a map feature extractor $g_{m}$, a sensor feature extractor $g_{s}$, and a fusing module $g_{f}$. The map and sensor features are fused, resulting in the control signal prediction
$(v,\omega)=g_{f}(g_{m}(M_{c},M_{o},M_{f}),g_{s}(S)).$ (7)
We consider three network architectures, a simple multilayer perceptron (MLP), a standard convolutional neural network (CNN), and our proposed scale-grouped convolutional neural network (SGCNN) that independently processes the different scales in the maps, see Fig. 5. The MLP is mainly used as a benchmark to evaluate the inductive priors in the CNN-based architectures. For the MLP, the feature extractors, $g_{m}$ and $g_{s}$, are identity functions that simply flatten the inputs, and the fusing module consists of three fully connected (FC) layers. The CNN-based architectures use convolutional layers in the map feature extractor followed by a single FC layer. In SGCNN, we group the maps in $M_{c}$, $M_{o}$ and $M_{f}$ by scale, as the pixel positions between different scales do not correspond to the same world coordinate. Each scale is convolved separately using grouped convolutions. This ensures that each convolution kernel is applied to grids where the spatial context is consistent across channels. The sensor feature extractor is a single FC layer, and the fusing module consists of three FC layers. More details can be found in Appendix A.
IV Transferring CPP Agents to the Real World
Training from scratch directly on the real robot is infeasible due to a slow convergence of the training process. In Section V-B, we find that the number of training iterations required is in the order of millions, which would translate to non-stop training for weeks to months in the real world. Meanwhile, it would require some form of manual interaction. Thus, we aim to transfer CPP agents trained in simulation, and fine-tune them on a real robot. In this section, we describe our proposed approach for sim-to-real transfer.
IV-A Learning CPP on the Real System
@@
Even if we will not train the agent from scratch in the real setting, we need to be able to fine-tune the transferred system by RL. In common RL-libraries, such as Stable-Baselines3 [45], Tianshou [46], and Spinning Up [47], the data collection and model updates are performed serially. While this is feasible in simulation and turn-based environments where the environment can be paused during model updates and time can be advanced instantly to the next state after action selection, it is not practical for real-time robotic applications where the agent is trained online.
Figure 5: Our proposed SGCNN architecture consists of convolution (CONV) and fully connected (FC) layers. The scales of the multi-scale maps are convolved separately as their spatial positions are not aligned in the grid. x3/x4 refer to the number of layers.
@@
In this case, we need to wait after action selection to observe the next state, while the execution remains idle. After the environment interaction, a batch is sampled from the replay buffer, gradients are computed, and the model is updated. However, this takes time, which is not negligible, especially on low-performance edge devices and embedded systems. During the gradient update step, the robot keeps following its previous action. As a result, the agent ends up in a state which deviates from the recorded one used for action selection. Following previous work [39, 14], we perform the model updates in parallel with the data collection, utilizing computational resources which would otherwise remain idle while waiting for the next state.
@@
The training process can be divided into four main computational steps, where the respective measured times are given in Table I:
@@
(1) Action selection: Given the current observation, we sample an action from the policy. This corresponds to a forward pass of the policy network.
@@
(2) State selection: After sampling an action, the control signals are sent to the robot platform, and the new state is observed. In simulation, this would occur instantly, while in the real world we need to let time pass to observe how the action affected the environment. After state selection, the reward is computed, and the action, previous state, new state, and reward are added to the replay buffer.
(3) Batch sampling: A training batch is sampled from the replay buffer.
@@
(4) Model update: Gradients are computed based on the training batch, and the model weights are updated. This is the most computationally intensive part.
TABLE I: Measured times of the RL-step on the real system. Timing for (2) is application dependent, and chosen based on the optimal training step size. The overhead effectively becomes part of (1) resulting in an action delay of 50 ms.
@@
Figure 6: During online training we perform environment interactions (top) and model training (bottom) in parallel. Either branch can be executed multiple times before weight synchronization.
Since the action selection and batch sampling were fast compared to the state selection and gradient computations, we only use two threads to reduce overhead and complexity. Our computational graph is shown in Fig. 6, which consists of an environment interaction thread (top) and a model update thread (bottom). Since both threads interact with the replay buffer, we use a mutex to avoid conflicts. Luckily, the environment interaction thread adds an entry at the end of the cycle, while the model update thread samples from it early, so the risk of the threads blocking each other is low.
Since the model update thread modifies the weight tensors during a large portion of the runtime, i.e. both during the backward pass to update the gradients and during the optimizer step to update the parameters, a mutex is not feasible as it would block execution too often. Instead, we keep a copy of the model, which is only accessed by the model update thread. The weights are synchronized when both threads have finished execution.
@@
Figure 7: A picture of the robot having covered an environment with four square obstacles in our semi-virtual setting.
IV-B Closing the Sim-to-Real Gap
With the approach in the previous section, we could directly move to a real setting in-the-wild. However, this would make training and fine-tuning cumbersome as manual interaction would be required to supervise the training process. In order to fully automatically (continue to) train the system, we propose to use a semi-virtual setup, see Fig. 7. In this setup, we use the real robot with its kinematics and dynamics, but simulate its lidar sensor and obstacles, and localize the robot using a positioning system. The environment, including obstacles and coverage, e.g. from mowing, is visualized by a projection onto the ground.
TABLE II: Physical dimensions for the three different CPP variations in simulation.
Omnidirectional Non-omnidirectional Lawn mowing
exploration exploration
Coverage radius 77 m $3.5$ m $0.15$ m
Agent radius $0.08$ m $0.15$ m $0.15$ m
Maximum linear velocity $0.5$ m/s $0.26$ m/s $0.26$ m/s
Maximum angular velocity 11 rad/s 11 rad/s 11 rad/s
Simulation step size $0.5$ s $0.5$ s $0.5$ s
Lidar rays 2020 $24$ $24$
Lidar range 77 m $3.5$ m $3.5$ m
Lidar field-of-view $360^{\circ}$ $^{\circ} 180180^{\circ}$
The main purpose of using such an environment is to evaluate, in a controlled manner, the ability of the learning algorithm to generalize to a real robot in a real-time setting. We can study how the agent adapts to real-world aspects, such as wheel slip, that occur here. Moreover, as we will see in Section V-B, the proposed RL approach is robust to sensing and localization noise, and can be extended to the case of unknown pose, e.g. by estimating the pose using an off-the-shelf SLAM method. Thus, using our semi-virtual environment, we can draw conclusions about how we expect our approach to generalize to a fully realistic setting.
@@ IV-B Closing the Sim-to-Real Gap
To further reduce the sim-to-real gap, we improve the fidelity of the simulator by taking into consideration the latencies induced by inertia and action delay. We measure the maximum linear acceleration, maximum angular acceleration, and action delay of the real system, and include these aspects in the simulated kinematics and dynamics.
To account for higher-order Markovian dynamics, we include information from previous time steps in the observation space. While the common approach is to stack several previous observations [39, 48], it is less feasible in our setting for multiple reasons. (1) Since the pose is embedded in the egocentric maps, rapid rotations significantly alter the observed state, making it difficult to learn the dynamics. (2) It significantly increases the model size and processing time, which are critical aspects for real-time applications. (3) It severely limits the replay buffer size, as the map observations are fairly large. Instead, we use a history of the previous actions, following previous work [39, 48]. This is lightweight and avoids the listed problems, and should be sufficient for learning the dynamics of the robot. Another option would be to use velocity estimates, although they can be highly noisy, in contrast to action observations. Using action observations, the agent can learn to estimate the velocity if necessary.
IV-C Optimal Strategy for Going Sim-to-Real
@@
    How does the model with fine-tuning on the real robot perform in comparison to (2) and depend on the number of training steps?
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d) (e) (f) (g) (h)
Figure 8: Examples of exploration maps (a-c), lawn mowing maps (d-f), and randomly generated maps (g-h).
V Experiments
In this section, we present our experimental results, both in terms of rigorous comparisons and ablations in simulation, as well as sim-to-real transfer to our semi-virtual setting.
V-A Implementation Details
Simulation training details. We first evaluate our RL approach in a simulated 2D environment. We utilize soft actor-critic learning [13], and train for 22-88M iterations with learning rate $2\cdot 10^{-5}$, batch size $256$, replay buffer size $5\cdot 10^{5}$, and discount factor $\gamma=0.99$. We evaluate our method on three settings: omnidirectional and non-omnidirectional exploration, as well as on the lawn mowing task. For physical dimensions in the different settings, see Table II. The training time for one agent varied between $25$ to $150$ hours on a T4 GPU and a 6226R CPU.